OpenStudy (sunnymadison):

Solve x2 + 9x + 9 = 0.

OpenStudy (xxdragoonxx):

x2 + 9x + 9 = 0 2+9x+9 = 11x+9=0 11x+9-9=0-9 11x = -9 11x/11 -9/11 x = -9/11

OpenStudy (kenzy1525):

where did you get the 11?

OpenStudy (adrimit):

is x2 = x^2?

jhonyy9 (jhonyy9):

x^2 +9x +9 = 0 solve it using discriminant D = b^2 -4ac so D= 81 -36 = 45 x_1,2 = (-9+/- sqrt45)/2 = (-9 +/- 3sqrt5)/2 hope this will help you

OpenStudy (mathmale):

Agree with jhonny. The discriminant (45) is positive; therefore, this quadratic equation has two real, unequal, roots. @sunnymadison: how would you check these results?